if p & q are positve single digit numbers find all possible solutions of 3p-4q=4
Answer by checkley77(12844)

3p-4q=4
3*4-4*2=4 (P=4,Q=2)
12-8=4
3*8-4*5=4 (P=8,Q=4)
24-20=4